. The activity of a source was measured as follows


Time/days 0 2 3 4 5


Activity 137 63 44 31 22


The background count was 7


i. Explain background count. How does it occur?


ii. Sketch the above graph and determine the half life of the source.

1
Expert's answer
2022-02-21T16:09:32-0500

"A=A_02^{-\frac tT,}"

"T=\frac t{\log_2\frac{A_0}A}=1.83^d."